OpenStudy (anonymous):

The difference between shelia and robbie's ages is 12. The sum of thier ages is 42. If shelia is the younger of the two. how old is robbie?

jimthompson5910 (jim_thompson5910):

let s = shelia's age r = robbie's age

jimthompson5910 (jim_thompson5910):

The difference between shelia and robbie's ages is 12 translates to r - s = 12 (note: shelia is the younger of the two) ----------------------------------------------- The sum of their ages is 42. translates to r+s = 42

jimthompson5910 (jim_thompson5910):

so you have this system r - s = 12 r+s = 42 add up the equations to get r - s = 12 r+s = 42 --------- 2r+0s = 54 so, 2r = 54 ---> r = ??? once you find r, use it to find s

OpenStudy (anonymous):

Thank you I am just

jimthompson5910 (jim_thompson5910):

yw